OpenStudy (czarluc):

2 sin(5x+4)=-square root 3

myininaya (myininaya):

Do you know how to solve 2sin(u)-sqrt(3) for u?

myininaya (myininaya):

there is suppose to be an equal sign there somewhere

myininaya (myininaya):

\[2 \sin(u)=-\sqrt{3}\]

OpenStudy (czarluc):

yes

myininaya (myininaya):

ok then solve that equation first then replace u with 5x+4 then solve for x

OpenStudy (czarluc):

why should you cancel out first the 5x+4?

myininaya (myininaya):

no one is canceling it

OpenStudy (czarluc):

uhm. i mean why should we forst get the value of 2sin=-square root 3?

myininaya (myininaya):

y=sin(x) and y=sin(5x+4) is basically the same graph except y=sin(x) is y=sin(5x+4)'s parent. y=sin(5x+4) is a transformation of the graph y=sin(x).

OpenStudy (czarluc):

oh. but should we give two cases of x since there are two values of u?

myininaya (myininaya):

Example: 2sin(x+1)=1 sin(x+1)=1/2 We don't even really have to replace x+1 with u. I just thought it would be easier for you. sin is 1/2 when x+1=pi/6+2npi or when x+1=5pi/6+2npi so we have x=pi/6+2npi-1 or when x=5pi/6+2npi-1

myininaya (myininaya):

y=sin(x+1) is just the graph of y=sin(x) moved to the left 1 unit

myininaya (myininaya):

That is why we got the answer answers as we would have gotten for sin(x) except -1

OpenStudy (czarluc):

why is there a 2npi?

myininaya (myininaya):

n*2pi 2pi is the circumference around a circle with radius 1 n represents how many circle rotations we make

OpenStudy (czarluc):

do we need to get the value of sin =1/2 first?

myininaya (myininaya):

sin(u)=1/2? if we are looking at my example yes

OpenStudy (czarluc):

last question.2cos^2(3x)+5cos(3x)-3=0 how about this?

myininaya (myininaya):

That is a quadratic in terms of cos(3x)

OpenStudy (czarluc):

just do the same? but many cases?

myininaya (myininaya):

solve the quadratic cos(3x) first \[2u^2+5u-3=0\]

myininaya (myininaya):

if it helps you can replace the cos(3x) with u first if it looks less nasty to you i did that above just now

OpenStudy (czarluc):

yeah I got it now thank you very much

OpenStudy (czarluc):

but if you use the general addition formula.... would the answer be the same?

myininaya (myininaya):

\[2\sin^2(x)+5\sin(x)-3=0 \\2\sin^2(x)+6\sin(x)-1\sin(x)-3=0 \\ 2\sin(x)[\sin(x)+3]-1[\sin(x)+3]=0 \\ (\sin(x)+3)(2\sin(x)-1)=0 \\ \text{ we have two equations \to solve } \sin(x)+3=0 \text{ or } 2\sin(x)-1=0 \\sin(x)=-3 \text{ or } \sin(x)=\frac{1}{2} \\ \text{ first off } \sin(x) \neq -3 \text{ since }\sin(x) \in [-1,1] \\ \sin(x)=\frac{1}{2} \text{ when } x=\frac{\pi}{6}+2npi \text{ or } x=\frac{5\pi}{6}+2npi\]

myininaya (myininaya):

That is just another example above having a quadratic in terms of a trig function

myininaya (myininaya):

what general addition formula?

OpenStudy (czarluc):

oh nevermind lol. just got it

OpenStudy (czarluc):

thanks!

myininaya (myininaya):

np